40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Entry  Mon Oct 14 17:01:43 2019, anchal, Notes, Laser, South Laser Switched Off! SouthLaserStopped.pdf
    Reply  Mon Oct 14 18:30:14 2019, anchal, Notes, PMC, South PMC Error Signal Oscillations SPMCMixerOutOscillations.pdf
       Reply  Tue Oct 15 18:54:16 2019, anchal, Notes, PMC, South PMC further problems 
          Reply  Wed Oct 16 16:56:43 2019, anchal, Notes, PMC, South PMC won't work with it's own Servo Card only 
             Reply  Wed Oct 16 17:33:08 2019, Koji, Notes, PMC, South PMC won't work with it's own Servo Card only 
                Reply  Wed Oct 16 18:46:58 2019, anchal, Notes, PMC, South PMC started working somehow :( 
                   Reply  Wed Oct 23 17:39:44 2019, anchal, Notes, PMC, South PMC started working somehow :( 
                      Reply  Fri Oct 25 16:50:12 2019, anchal, Notes, PMC, South PMC problem debugging efforts 
                         Reply  Tue Oct 29 19:10:59 2019, anchal, Notes, PMC, South PMC problem debugging efforts SPMC_Diagnostics.pdf
                            Reply  Wed Oct 30 15:15:28 2019, anchal, Notes, PMC, Replacing U5A AD602 didn't solve the issue. 
                               Reply  Wed Oct 30 18:35:31 2019, anchal, Notes, PMC, Issue closed 
                                  Reply  Mon Nov 4 12:15:01 2019, anchal, Notes, PMC, Changing the autolocking method 
Message ID: 2464     Entry time: Wed Oct 23 17:39:44 2019     In reply to: 2458     Reply to this: 2466
Author: anchal 
Type: Notes 
Category: PMC 
Subject: South PMC started working somehow :( 

The same thing happened again. This time, not just with the SPMC actuation voltage, but the South Laser slow voltage control was also unresponsive. However, I am not very sure about the latter. This was resolved once the restarted the whole lab. This narrows down the problem to following possibilities:

  • One or more channels in an acromag card become unresponsive. It could be the 10.0.1.49 acromag used for South PMC Controls.
  • Some device in TCS or QIL lab is interfering by using an IP address reserved for docker services on which the EPICS channels are hosted and python scripts are run.

These still don't explain CTN:2456. Again, since this is an irreproducible error, I will just have to wait for it to happen to gather more clues. Right now, everything is fine and beatnote is traveling towards set frequency.

 

ELOG V3.1.3-